OpenStudy (anonymous):

Write the equation of a line in slope-intercept form passing through (3, 5) and (2, 9)

OpenStudy (anonymous):

is it in y=mx+c??

OpenStudy (anonymous):

m=( 9-5)/ (2-3) = -4 y=mx+c 5 =-4(3) + c c= 17 y= -4x +17